A particle of mass 'm' moves in x-y plane so that its position vector r = xi + yj with a parametric equation x= a cos wt, y= b sin wt, where a&b are constants, a>b and w is a positive constant. 1. Show that the force field is conservative. 2. Find the potential energy at point a&b. 3. Find the work done by the force in moving the particle from point a-b. 4. Find the total energy of the particle.


1
Expert's answer
2021-06-08T09:34:57-0400

Given,

Mass of the particle = m

Position vector of the particle "(r) = x\hat{i} + y\hat{j}"

"x=a\cos(\omega t)"

"y=b\sin(\omega t)"

"a>b" and "\omega>0"

Now, we can write the position vector as,

"\overrightarrow{r}=a\cos(\omega t)\hat{i}+b\sin (\omega t)\hat{j}"

Now, "\overrightarrow{v}=\frac{dr}{dt}=-a\omega \sin(\omega t)\hat{i}+b\omega\cos(\omega t)\hat{j}"

Acceleration of the particle, "(\overrightarrow{a}) = \frac{dv}{dt}"

"\overrightarrow{a}=-a\omega^2 \cos(\omega)\hat{i}-b\omega^2\sin(\omega t)\hat{j}"

i) "F= m\overrightarrow{a}"

"F=-m\omega^2(acos(\omega t)\hat{i}+b\sin(\omega t)\hat{j})"

"=-m\omega^2(\overrightarrow{r})"

So, the force is depending on the angle.

ii) Velocity will be maximum for the case a, when "\omega t=0"

In that case "V=a"

Velocity will be maximum, when "\omega t = \pi/2"

"V=b"

Hence, the required potential energy "PE = \frac{1}{2}ma^2\omega^2" and "PE=\frac{1}{2}mb^2\omega^2"

iii)